OpenStudy (anonymous):

Maria solved the equation below and got x = –4. Which of the following choices verifies that her solution is correct? 5x + 22 = x + 6 A. 5(–4) + 22 = (–4) + 6 –20 + 22 = –2 –2 = –2 B. 5(–31) + 22 = (–10) + 6 –26 + 22 = –4 –4 = –4 C. 5(–4) + 22 = (–4) + 6 5(18) = (–4) +6 24 = 24 D. 5(–4) + 22 = (–4) + 6 –20 + 22 = 2 2 = 2

OpenStudy (dmb1215):

Not B

OpenStudy (anonymous):

No, but you have to plug in -4 into it.

OpenStudy (anonymous):

So plug it in and see what you get..

OpenStudy (anonymous):

is it a

OpenStudy (dmb1215):

no its D

OpenStudy (anonymous):

5(–4) + 22 = (–4) + 6 -20 + 22 = -4 + 6 2 = 2

OpenStudy (anonymous):

Yeah, the answer is D.

OpenStudy (anonymous):

thanks again guys

OpenStudy (dmb1215):

anytime